OpenStudy (anonymous):

2= sec (theta) + tan (theta) how to solve this?can anybody help?

OpenStudy (anonymous):

solve mean what?

OpenStudy (anonymous):

u want to prove it?

OpenStudy (anonymous):

find theta ,the ans is 0.644 but i could't get it

OpenStudy (kira_yamato):

Will leave the rest to you...

OpenStudy (anonymous):

wao.thanks!! im new but i've alrdy like it here!

OpenStudy (anonymous):

Sec(theeta)-Tan (theta)=half because sec square theta - tan square theta=1 hence sec theta=5/4 and tan theta=3/4

OpenStudy (anonymous):

how do u get there???

OpenStudy (anonymous):

thats incorrect

OpenStudy (jamesj):

yes, what vvvv has written is incorrect. But what Kira has written is very good ... use that.